logx(1/16)=-4
the log x, the x is suppose to be under the log.
here are the choices:
a. 2
b. 1/4
c. 4
d. 1/2

Answer by nerdybill(7384)

.
logx(1/16)=-4
1/16 = x^(-4)
1/16 = 1/x^4
Cross multiplying:
x^4 = 16
x = fourth root (6)
x = 2
.
Answer: a